Zinc reacts with hydrochloric acid according to the reaction equation


Zn(s)+2HCl(aq)⟶ZnCl2(aq)+H2(g)


How many milliliters of 3.50 M HCl(aq) are required to react with 2.35 g Zn(s)?


1
Expert's answer
2022-04-06T22:48:05-0400

Zn(s) + 2HCl(aq) ⟶ ZnCl2(aq) + H2(g)

n(Zn) = m(Zn)/Mr(Zn) = 2.35 g / 65.38 g/mol = 0.0359 mol

1 mol Zn - 2 mol HCl

0.0359 mol - x mol HCl

x = 0.0359 × 2 = 0.072 mol

V(sol) = n(HCl)/C(sol) = 0.072 mol / 3.50 mol/L = 0.02057 L = 20.57 mL